OpenStudy (anonymous):

If f(x) = |x – 5| + 2, find f(3).

OpenStudy (anonymous):

plug in 3 into the equation and solve.

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

Put 3 in place of x ....like this: f(3)=3-5+2=-2+2=0

OpenStudy (anonymous):

\(|x-5|=x-5\) \(|x-5|=5-x\)

OpenStudy (anonymous):

thank you got it

OpenStudy (neer2890):

it should be 4 because here we are working for mod function and |3-5|+2=|-2|+2=2+2=4

OpenStudy (neer2890):

how can you say that x=0 @Sourav12084

OpenStudy (neer2890):

f(x)=0